OpenStudy (anonymous):

Can someone help me better understand this information. Explain how to factor the following trinomials forms: a. x^2 + bx + c and b. ax^2 + bx + c.

OpenStudy (vishweshshrimali5):

lets take an example of each \[x^2 + 2x + 1\] \[= x^2 + x + x + 1\] \[= x(x+ 1) + 1(x+1)\] \[=(x+1)(x+1)\] \[=(x+1)^2\]

OpenStudy (anonymous):

Breaking it down that way makes it easier for me to understand it. I believe the variables throw me off.

OpenStudy (vishweshshrimali5):

practice will make u perfect

OpenStudy (vishweshshrimali5):

\[4x^2 + 12 x + 9\] \[4x^2 + 6x + 6x + 9\] \[2x(2x + 3) + 3(2x + 3)\] \[(2x+3)(2x+3)\] \[(2x+3)^2\]

OpenStudy (vishweshshrimali5):

clear ?

OpenStudy (anonymous):

Much better thank you

OpenStudy (vishweshshrimali5):

never mind !